# Nagaoka Ferromagnetism in $3\times 3$ Arrays and Beyond

### v2

@@ A.7 Scaling the transition from nNF to NF for $N\times N$ arrays
The transition from NF to nNF as $U$ decreases from the infinite $U$ limit occurs when the energy shift from the exchange coupling to the states with a doubly-occupied site closes the gap between the excited state and the fully polarized ground state. Using $\mathcal{H}_{sc}$, the second-order exchange $-(t/U)^{2}$ closes the gap which scales between $|t|/UN^{3}$ and $|t|/UN^{4}$. As a consequence, the scaling of $|t|/U$ at the transition should be between $1/N^{4}$ and $1/N^{3}$. This is consistent with our calculations for different $N$ which show a $1/N^{3.5}$ scaling.
@@ B.2 Two-hole IF in plaquette sharing corners
5 corner-sharing sq [Uncaptioned image] IF 14 $0.0002$ 2-hole IF
@@
B.3 Robustness to extended Coulomb interaction

### v3

@@ A.7 Scaling the transition from nNF to NF for $N\times N$ arrays
The transition from NF to nNF as $U$ decreases from the infinite $U$ limit occurs when the energy shift from the exchange coupling to the states with a doubly-occupied site closes the gap between the excited state and the fully polarized ground state. Using $\mathcal{H}_{sc}$, the second-order exchange $-(t/U)^{2}$ closes the gap which scales between $|t|/UN^{3}$ and $|t|/UN^{4}$. As a consequence, the scaling of $|t|/U$ at the transition should be between $1/N^{4}$ and $1/N^{3}$. This is consistent with our calculations for different $N$ which show a $1/N^{4}$ scaling. When we include all values of $N$ to determine the scaling, the scaling is closer to $1/N^{3.5}$ scaling. The scaling for the larger $N$ should depend of the state splitting near the band edge and should have a scaling closer to $1/N^{4}$, consistent with the scaling we get for larger $N$ in the computational results. When the smaller $N$ are included in the scaling, then the small $N$ results will involve states further from the band edge which provide a $1/N^{3}$ splitting and scaling. The scaling using all $N$ should lie between $1/N^{3}$ and $1/N^{4}$, consistent with the $1/N^{3.5}$ scaling seen in the calculated results when all $N$ are considered.
@@ B.2 Two-hole IF in plaquette sharing corners
5 corner-sharing sq [Uncaptioned image] IF 14 $0.00025$ 2-hole IF
@@
B.3 Adjacent Loops
For small arrays made of adjacent loops, there are special fillings which give rise to saturated ferromagnetism. The fillings are related to the number of loops present in the system. For two adjacent loops (calculated by ED for up to two 14-site loops), with the two loops sharing one hopping and two sites, there exists 5-electron ferromagnetism when the two loops are about the same size. The dependence of the transition point on total system size N is shown below. Similar situation also happens for three adjacent loops with 7-electron ferromagnetism for certain geometries. How the loop number and special fillings are related will require further investigation.
Figure 12: Two adjacent loops. Transition point $|t/U|$ versus total system size $N$
B.4 Robustness to extended Coulomb interaction